OpenStudy (anonymous):

f(x)=3csc(pi/4x+pi/8) I need to know a. the period of this function, b. the sequence of transformations from g(x)= csc(x) -> F(x) = 3 csc(pi/4x+pi/8)

OpenStudy (mertsj):

Same as before. Factor out the pi/4 first.

OpenStudy (mertsj):

The period can be found with this formula \[Coefficient` of `x = \frac{2 \pi}{period}\]

OpenStudy (anonymous):

ok so its starts with the vert stretch from 1 -> 3, then a horizontal stretch by a factor of \[\frac{ \pi }{ 4}\]

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

then a shift to the left of......and I'm stuck. i want to say \[\left(\begin{matrix}\pi \\ 2\end{matrix}\right) -\left(\begin{matrix}\pi \\ 42\end{matrix}\right)\]

OpenStudy (mertsj):

Shift to left of pi/2

OpenStudy (anonymous):

is that from pi/4* pi/8?

OpenStudy (mertsj):